Kathy had a party. Everyone at the party shook hands with everyone else.
How many people were at the party if there 78 handshakes?

Answer by askmemath(368)

Have you studied Permutations and Combinations?
IF there were only 2 people,then only 1 handshake wld have occured
If 3 ppl, then 3 hand shakes
If 4 people then 6
If n people then its nC2 or n(n-1)/2
So if 78 handshakes = n(n-1)/2 = 78
n(n-1) = 78x2
= 39 x2 x2
= 13x3x2x2
= 13x12
So 12people were at the party